# Benchmarking stabilized and self-stabilized $p$-virtual element methods with variable coefficients

### v1

@@ 8.1.2 Example 2: trigonometric coefficient
$\mathcal{A}\left(x,y\right)=\begin{bmatrix}\cos^{4}\left(\pi x\right)+1&\cos^{2}\left(\pi x\right)\sin^{2}\left(\pi y\right)\\ \cos^{2}\left(\pi x\right)\sin^{2}\left(\pi y\right)&sin^{4}\left(\pi y\right)+3\end{bmatrix}$. (77)
@@ 8.2 Linear elasticity problem
Lastly, the results for the linear elasticity are presented. The same exact solution, mesh types and order $p$ of Section 6.2 are considered. By following the idea outlined in [reddy2019virtual], the projection operators for standard VEM are computed by considering a piece-wise constant approximation of the coefficient. In each element, $\mathbb{C}$ is approximated by its average at the integration points. The error is measured as in Eq. (63) and the analysis is conducted with the stabilized VEM $\bm{S}^{E}_{3}$ ($\tau=0.5$) and the self-stabilized formulation $\mathrm{V1}$ ($\mathrm{tol}=\mathrm{tol}_{0}$), which has not been previously investigated.
@@
For the quadrilateral mesh, results are shown in Figure 21. The analysis conducted with the newly developed formulation retains optimal accuracy. This is true both for the $L^{2}$ and for the energy error. Also, the self-stabilized $\mathrm{V1}$ and the approach presented in [daveiga2016virtual], extended to the present problem, provide an adequate description of the solution.
@@
The results obtained with the Bézier-edge mesh are presented in Figure 22 and are similar to the previous case. The approach discussed in [daveiga2016virtual] features optimal accuracy up to order $p=9$ with a slight decrease for $p\geq 10$. Therefore, to determine whether this accuracy loss is an isolated behavior or not, this analysis is conducted up to order $p=12$. It is clear that for $p\geq 10$ the $\Pi^{0}_{p-1}$ stabilized VEM experiences a flattening of the error curve for this mesh type. On the contrary, the $\mathrm{VC}$-VEM simulations maintain the accuracy both in the $L^{2}$ and in the energy error. The same happens for the self-stabilized $\mathrm{V1}$, despite a slight error increase is observed for low $p$. As in the previous tests, the standard stabilized VEM features larger errors as the order $p$ increases.
@@ 9 Conclusions
The second part of the paper focused on the case of variable coefficients, with the introduction of a new virtual element approach, denoted by $\mathrm{VC}$-VEM. $\mathrm{VC}$-VEM consists in modifying the discrete space so that the general coefficient can be taken into account when computing the polynomial projection. The newly introduced formulation was compared with the standard approaches and provided promising results. The comparison can be summarized by the following highlights.
@@
    As already observed in [daveiga2016virtual], the stabilized VEM constructed with the operator $\Pi^{\nabla}_{p}$ has a strong loss of accuracy for $p\geq 3$.
@@
    The stabilized VEM constructed by the operator $\Pi^{0}_{p-1}$ provides optimal results for polynomial coefficients, while its accuracy deteriorates in the presence of more general coefficients.
@@
Owing to the obtained results, future developments will apply the present formulation to the analysis of variable stiffness panels in the nonlinear regime.

### v2

@@ 8.1.2 Example 2: trigonometric coefficient
$\mathcal{A}\left(x,y\right)=\begin{bmatrix}\cos^{4}\left(\pi x\right)+1&\cos^{2}\left(\pi x\right)\sin^{2}\left(\pi y\right)\\ \cos^{2}\left(\pi x\right)\sin^{2}\left(\pi y\right)&\sin^{4}\left(\pi y\right)+3\end{bmatrix}$. (77)
@@ 8.2 Linear elasticity problem
Lastly, the results for the linear elasticity are presented. The same exact solution, mesh types and order $p$ of Section 6.2 are considered. By following the idea outlined in [32], the projection operators for standard VEM are computed by considering a piece-wise constant approximation of the coefficient. In each element, $\mathbb{C}$ is approximated by its average at the integration points. The error is measured as in Eq. (63) and the analysis is conducted with the stabilized VEM $\bm{S}^{E}_{3}$ ($\tau=0.5$) and the self-stabilized formulation $\mathrm{V1}$ ($\mathrm{tol}=\mathrm{tol}_{0}$), which has not been previously investigated.
@@
For the quadrilateral mesh, results are shown in Figure 21. The analysis conducted with the newly developed formulation retains optimal accuracy. This is true both for the $L^{2}$ and for the energy error. Also, the self-stabilized $\mathrm{V1}$ and the approach presented in [7], extended to the present problem, provide an adequate description of the solution.
@@
The results obtained with the Bézier-edge mesh are presented in Figure 22 and are similar to the previous case. The approach discussed in [7] features optimal accuracy up to order $p=9$ with a slight decrease for $p\geq 10$. Therefore, to determine whether this accuracy loss is an isolated behavior or not, this analysis is conducted up to order $p=12$. It is clear that for $p\geq 10$ the $\Pi^{0}_{p-1}$ stabilized VEM and the self-stabilized $\mathrm{V1}$ experience a flattening of the error curve for this mesh type. On the contrary, the $\mathrm{VC}$-VEM simulations maintain the accuracy both in the $L^{2}$ and in the energy error. As in the previous tests, the standard stabilized VEM features larger errors as the order $p$ increases.
@@ 9 Conclusions
  • 3.
    The stabilized VEM formulations retain at least the same accuracy as the self-stabilized ones, while preserving better conditioning and lower computational cost.
The second part of the paper focused on the case of variable coefficients, with the introduction of a new virtual element approach, denoted by $\mathrm{VC}$-VEM. It consists in defining a new polynomial projection which combines the $L^{2}$ projector and the features of the coefficient in the framework of standard VEM spaces. The newly introduced formulation was compared with the standard approaches and provided promising results. The comparison can be summarized by the following highlights.
@@
    As already observed in [7], the stabilized VEM constructed with the operator $\Pi^{\nabla}_{p}$ has a strong loss of accuracy for $p\geq 3$.
@@
    The stabilized VEM constructed by the operator $\Pi^{0}_{p-1}$ provides optimal results for polynomial coefficients, while a slight accuracy deterioration is observed at high approximation orders ($p\geq 6$) in the case of the trigonometric coefficient.
@@
Building on the obtained results, future developments will address the extension of the present formulation to geometrically nonlinear problems in elasticity.